OpenStudy (anonymous):

sketch the graph, y=-3x+2 give the solution set

OpenStudy (anonymous):



OpenStudy (anonymous):

I can graph it but I don't know the solution set

OpenStudy (anonymous):

solution set from what?

OpenStudy (anonymous):

if you mean the solution set with y=x, then y=-3x+2 and y=x and x=1/2,y=1/2